If one side of a square is tripled and another side is increased by 4cm,the area of the resulting rectangle is 180cm^2.What is the length of the side of the original square?
Answer by mananth(16946)

let side of square be x
3x & x+4 are the new dimensions
Area = L * W
3x(x+4) = 180
3x^2+12x=180
3x^2+12x-180=0
/3
x^2+4x-60=0
x^2+10x-4x-60=0
x(x+10)-4(x+10)=0
(x+10)(x-4)=0
x=-10 OR 4
x cannot be negative
therefore the side of square is 4cm